OpenStudy (anonymous):

i need help with integration of rational functions by partial fractions, the question is x+4/(x+1)(x+1)^2 please help

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{x+4}{(x+1)(x+1)^2}~dx}\) this?

OpenStudy (solomonzelman):

for this one, you don't need partial fractions. \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{x+4}{(x+1)(x+1)^2}~dx}\) \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{x+4}{(x+1)^3}~dx}\) u=x+1, du=dx, x=u-1 ---> x+4=u-3 \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{u-3}{u^3}~du}\) break it up, and integrate using the power rule, then sub the x+1 for u back as you are done.

OpenStudy (solomonzelman):

on top supposed to be u+3, not u-3. it is a typo, sorry.

OpenStudy (anonymous):

with out the integral

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\frac{x+4}{(x+1)(x+1)^2}}\) you want partial fractions for this?

OpenStudy (anonymous):

yeah the book has that

OpenStudy (anonymous):

shoot its x+4/(x+1)^2 im sorry

OpenStudy (solomonzelman):

its ok \(\large\color{black}{ \displaystyle \frac{x+4}{(x+1)^2}= \frac{A}{x+1}+\frac{B}{(x+1)^2}}\)

OpenStudy (solomonzelman):

this is the setup.

OpenStudy (anonymous):

yeah that is what i have

OpenStudy (solomonzelman):

solve for A and B, and re-write it

OpenStudy (solomonzelman):

at first multiply every fractions time (x+1)^2

OpenStudy (solomonzelman):

you will get \(\large\color{black}{ \displaystyle x+4=A(x+1)+B}\)

OpenStudy (anonymous):

so then the x value would be -1

OpenStudy (solomonzelman):

yes, you can solve for B like that.

OpenStudy (solomonzelman):

B=?

OpenStudy (anonymous):

3

OpenStudy (solomonzelman):

yes, B=3

OpenStudy (solomonzelman):

now, you know that B=3, and lets choose an alternate x-value (say, x=1) \(\large\color{black}{ \displaystyle x+4=A(x+1)+B}\) \(\large\color{black}{ \displaystyle 1+4=A(1+1)+3}\)

OpenStudy (solomonzelman):

solve for A.

OpenStudy (solomonzelman):

or choosing x=0 is even better

OpenStudy (anonymous):

oh ok there we go i wasnt putting 3 in

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 0+4=A(0+1)+3}\)

OpenStudy (solomonzelman):

yes, you got to put that 3 for B, or else you have no way to find A:)

OpenStudy (solomonzelman):

A=?

OpenStudy (anonymous):

1

OpenStudy (solomonzelman):

yes A=1

OpenStudy (solomonzelman):

And let me verify it

OpenStudy (solomonzelman):

here is the \(\scr\color{blue}{\href{ http://www.wolframalpha.com/input/?i=%28x%2B4%29%2F%28%28x%2B1%29%5E2%29+partial+fractions }{verification }}\).

OpenStudy (solomonzelman):

oh, forgot another / here is the \(\scr\color{blue}{\href{http:///www.wolframalpha.com/input/?i=%28x%2B4%29%2F%28%28x%2B1%29%5E2%29+partial+fractions}{verification }}\).

OpenStudy (solomonzelman):

was supposed to have 3 times \ before wolfram, but just in case you now know how to make hyperlinks. ``` \(\scr\color{blue}{\href{http:///www.wolframalpha.com/input/?i=%28x%2B4%29%2F%28%28x%2B1%29%5E2%29+partial+fractions}{verification }}\). ```

OpenStudy (solomonzelman):

bye

OpenStudy (anonymous):

thanks it helped